OpenStudy (anonymous):

Just a quick alg question, its stumped me. Find the value of 3 over the sqrt of 3

OpenStudy (anonymous):

\[\frac{3}{\sqrt{3}} \implies \frac{3}{\sqrt{3}} \times \frac{\sqrt{3}}{\sqrt{3}} \implies \frac{3\sqrt{3}}{3} \implies ?\]

OpenStudy (anonymous):

oh....i understand now

OpenStudy (anonymous):

Yeah. It's just rationalizing the denominator.

OpenStudy (anonymous):

It would just be the sqrt(3)...Thank you very much

OpenStudy (anonymous):

np :)